# Flow Matching and Diffusion Models via PointNet for Generating Fluid Fields on Irregular Geometries

### v1

@@ Abstract
We present two novel generative geometric deep learning frameworks, termed Flow Matching PointNet and Diffusion PointNet, for predicting fluid flow variables on irregular geometries by incorporating PointNet into flow matching and diffusion models, respectively. In these frameworks, a reverse generative process reconstructs physical fields from standard Gaussian noise conditioned on unseen geometries. The proposed approaches operate directly on point-cloud representations of computational domains (e.g., grid vertices of finite-volume meshes) and therefore avoid the limitations of pixelation used to project geometries onto uniform lattices. In contrast to graph neural network–based diffusion models, Flow Matching PointNet and Diffusion PointNet do not exhibit high-frequency noise artifacts in the predicted fields. Moreover, unlike such approaches, which require auxiliary intermediate networks to condition geometry, the proposed frameworks rely solely on PointNet, resulting in a simple and unified architecture. The performance of the proposed frameworks is evaluated on steady incompressible flow past a cylinder, using a geometric dataset constructed by varying the cylinder’s cross-sectional shape and orientation across samples. The results demonstrate that Flow Matching PointNet and Diffusion PointNet achieve more accurate predictions of velocity and pressure fields, as well as lift and drag forces, and exhibit greater robustness to incomplete geometries compared to a vanilla PointNet with the same number of trainable parameters.
@@ 1 Introduction and motivation
Flow matching [1], diffusion models [2], and their advanced variants [3, 4, 5] are generative deep-learning frameworks that have demonstrated success in computer graphics and computer vision, particularly for image, video generation [6], outperforming earlier generative approaches such as generative adversarial networks (GANs) [7, 8] and variational autoencoders (VAEs) [9]. At a high level, in supervised flow matching [1] and diffusion [2] frameworks, a forward process adds noise to the training data, while an inverse process predicts the clean fields. Several formulations of flow matching and diffusion models operate in a conditional setting, where generation is guided by auxiliary inputs such as textual prompts [4, 10] or low-resolution images [11, 12]. Motivated by their success in computer graphics, these models have been extended to computational physics, enabling a range of applications including predictions conditioned on material properties [13, 14, 15], high-fidelity field reconstruction from low-resolution data [16, 17], and prediction of physical fields conditioned on the geometry of the computational domain [13, 18, 19]. In this article, we focus on the latter case.
@@
Flow matching and diffusion models conditioned on the geometry of the computational domain fall within the category of geometric deep learning. These approaches are critical in computational physics, as they accelerate shape-optimization studies that require exploration of high-dimensional geometric parameter spaces [20]. Generative flow matching and diffusion models conditioned on geometry have been shown to outperform regression-based surrogate models for predicting physical fields (see e.g., Refs. [21, 22, 23]), such as convolutional neural networks (CNNs) [24], Fourier neural operators (FNOs) [25], and deep operator neural networks (DeepONets) [26], which directly map geometry to physical fields. In particular, traditional CNNs [24], FNOs [25], and DeepONets [26] produce a single output for a given geometry, whereas flow matching and diffusion models can generate multiple realizations for the same geometry, thereby quantifying model uncertainty. Moreover, diffusion models have demonstrated superior performance compared to GANs and VAEs in predicting physical field variables (see, e.g., Refs. [21, 27]). In geometric deep learning using flow matching and diffusion models, two steps are particularly important. First, how to represent irregular and complex geometries; and second, how to condition flow matching and diffusion models on geometry from a computer science perspective.
@@
For the representation of irregular geometries, two common approaches are used. The first is a pixelation-based strategy, in which the geometry is projected onto uniform Cartesian grids. In this setting, Q. Liu and N. Thuerey [13] proposed a diffusion model conditioned on the geometry of airfoils to predict velocity and pressure fields. In their work, the airfoil geometry was represented using the pixelation approach, and a U-Net architecture [28, 2] was employed as the backbone neural network. The authors [13] performed geometric conditioning by concatenating the binary image of the airfoil with the velocity and pressure fields as the input to the U-Net, while noise was added only to the velocity and pressure fields during training (see Fig. 3 of Ref. [13]). A similar approach was taken by Zheng et al. [18], where the authors employed an encoder that maps a two-dimensional binary image of the geometry to a latent feature representation, which was then used in a diffusion transformer to condition the diffusion model on the geometry (see Fig. 1 of Ref. [18]). Similarly, J. Hu et al. [19] proposed a diffusion model conditioned on the geometry of obstacles to predict incompressible flow fields. The authors [19] conditioned the diffusion model on geometry by encoding the obstacle as a binary mask through a U-Net–style encoder and injecting this geometry representation into the diffusion process via cross-attention blocks (see Fig. 1 of Ref. [19]). Their results demonstrated that the proposed diffusion model outperformed CNNs and VAEs [19].
@@
The Flow Matching PointNet and Diffusion PointNet frameworks consist of a forward process and an inverse denoising process. In Diffusion PointNet, the forward process incrementally corrupts ground-truth velocity and pressure fields in fluid flows by adding standard Gaussian noise. PointNet is then trained to predict the noise added at each diffusion step [2]. This prediction is conditioned on the spatial coordinates of the point cloud, which represent the domain geometry, as well as on a sinusoidal time embedding. The conditioning is implemented by concatenating the spatial-coordinate vectors and the sinusoidal-embedding vector with the corrupted velocity and pressure fields. In the inverse process, generation begins from samples drawn from a standard Gaussian noise distribution and is conditioned on the spatial coordinates of the point cloud and the sinusoidal embedding. During the inverse procedure, the trained PointNet predicts the noise at each stage and progressively reconstructs the velocity and pressure fields in a manner consistent with the imposed conditioning. In the Flow Matching PointNet model, PointNet is trained to predict a target vector field that corresponds to the temporal derivative of a continuously perturbed field with respect to an auxiliary time variable [1]. In the inverse process, this learned vector field is used to numerically integrate an ordinary differential equation, progressively reconstructing the velocity and pressure fields from an initial standard Gaussian noise distribution [1], using the same conditioning mechanism as in Diffusion PointNet. To investigate the performance of Flow Matching PointNet and Diffusion PointNet, the classical fluid dynamics problem of incompressible flow past a cylinder is considered, in which the cross-sectional shape of the cylinder varies across the dataset.
Note that combinations of diffusion models and point cloud neural networks in the computer graphics literature have been used in which, during the forward process, noise is added to the spatial coordinates of shapes and, during the inverse process, new three-dimensional geometries are generated [41, 48]. In contrast, in the present study, noise is added to and removed from the physical fields, while the spatial coordinates remain clean and unchanged. Hence, the proposed Flow Matching PointNet and Diffusion PointNet can be extended for use in part and semantic segmentation of three-dimensional point clouds in computer graphics.
@@ 2 Data preparation
Fig. 2: Architecture of Flow Matching PointNet, Diffusion PointNet, and Baseline PointNet. Labels in the form $(\mathcal{B}_{1},\mathcal{B}_{2})$ and $(\mathcal{B}_{1},\mathcal{B}_{2},\mathcal{B}_{3})$ are used to denote shared MLPs, as illustrated in Sect. 3.4. $n_{\mathrm{CFD}}$ denotes the number of CFD variables and $N$ represents the number of points in the point clouds. In the Flow Matching PointNet architecture, the subscript $\tau$ is used to indicate $\boldsymbol{s}_{\tau}$ and $\boldsymbol{y}_{\tau}$ (see Sect. 3.1). In the Diffusion PointNet architecture, the subscript $t$ is used to show $\boldsymbol{s}_{t}$ and $\boldsymbol{y}_{t}$ (see Sect. 3.2). In Flow Matching PointNet and Diffusion PointNet, the outputs are, respectively, the prediction of the target field ($\boldsymbol{f}_{\text{target}}$) and the noisy field ($\boldsymbol{y}_{\text{noisy}}$), given the polluted velocity and pressure fields at steps $\tau$ and $t$, and conditioned on the geometry of the physical domain represented as a point cloud.
@@ 3.1 Generative Flow Matching PointNet
After training, the Flow Matching PointNet generates the velocity and pressure fields conditioned on a geometry from the test set. Given $\boldsymbol{x}$ as the point cloud representing the target geometry from the test set, we generate three independent standard Gaussian fields representing $\boldsymbol{y}_{\text{noisy}}$ and integrate Eq. 12 from $\tau=1$ to $\tau=0$ using the explicit Euler method with $N_{\tau}$ steps, such that $\Delta\tau=1/N_{\tau}$. At each step,
@@ 3.2 Generative Diffusion PointNet
Similar to Sect. 3.1, let $\boldsymbol{y}_{\text{clean}}\in\mathbb{R}^{N\times n_{\text{CFD}}}$ represent the clean field, where $N$ and $n_{\text{CFD}}$ are defined as before and set to 1024 and 3, respectively, since our objective is to predict the two-dimensional velocity and pressure variables $u^{\prime}$, $v^{\prime}$, and $p^{\prime}$. Figure 3 illustrates the schematic of the forward and inverse processes in Diffusion PointNet. In the forward diffusion process, standard Gaussian noise is incrementally added to $\boldsymbol{y}_{\text{clean}}$ over $T$ discrete steps using a noise scheduler, producing a sequence of noisy fields $\{\boldsymbol{y}_{t}\}_{t=1}^{T}$, where $\boldsymbol{y}_{t}\in\mathbb{R}^{N\times n_{\text{CFD}}}$. In this study, we use a cosine noise scheduler defined as
@@
where $\epsilon_{u_{i}}$, $\epsilon_{v_{i}}$, and $\epsilon_{p_{i}}$ denote the ground-truth standard Gaussian noise, as similarly defined in Sect. 3.1, and $\tilde{\epsilon}_{u_{i}}$, $\tilde{\epsilon}_{v_{i}}$, and $\tilde{\epsilon}_{p_{i}}$ denote the corresponding noise components predicted by the Diffusion PointNet. Diffusion PointNet is optimized with the Adam algorithm [61] using a constant learning rate of 0.001 and a batch size of 256. For each training iteration, a random step $t$ is sampled independently for every point cloud in the training batch, and three independent standard Gaussian noises are generated, corresponding to the $u^{\prime}$, $v^{\prime}$, and $p^{\prime}$ variables, to construct the tensor $\boldsymbol{\epsilon}$ presented in Eq. 19. The corresponding noisy tensor $\boldsymbol{y}_{t}$ is then computed using Eq. 19.
Once training is complete, the Diffusion PointNet generates the corresponding velocity and pressure variables for a geometry selected from the test set by simulating the reverse diffusion process, as shown in Fig. 3. For a point cloud $\boldsymbol{x}$ representing a target geometry, the reconstruction begins by generating three independent standard Gaussian noise fields to construct the tensor $\boldsymbol{y}_{T}$. The fields are then iteratively denoised from $t=T$ to $t=0$ using
@@ 3.3 Baseline PointNet
Now that the Flow Matching PointNet and Diffusion PointNet architectures have been described, the baseline PointNet can be introduced in a straightforward manner. The baseline PointNet for predicting fluid flow variables was first proposed by Kashefi et al. [29] and has since been adopted in several subsequent frameworks (see e.g., Refs. [63, 30, 64]). The version considered in this work is identical to that originally presented by Kashefi et al [29]. In the baseline PointNet, the input consists solely of the spatial coordinates of the point cloud representing the geometry of the domain, and the output corresponds to the physical fields of interest, as shown in the bottom panel of Fig. 2. Following the notation introduced in Sect. 3.1, the input is given by $\boldsymbol{x}\in\mathbb{R}^{N\times d}$, where in the present study $d=2$, corresponding to a two–dimensional problem with coordinates $(x^{\prime},y^{\prime})$, and $N=1024$, which denotes the number of points per point cloud. Consequently, the input tensor is $\mathcal{X}\in\mathbb{R}^{N\times d}$. The output consists of the velocity components and pressure field $(u^{\prime},v^{\prime},p^{\prime})$, forming an output tensor $\mathcal{Y}\in\mathbb{R}^{N\times n_{\mathrm{CFD}}}$. Hence, PointNet learns a direct mapping from $\mathcal{X}$ to $\mathcal{Y}$. The loss function is the mean squared error and can be written as
@@ 3.4 PointNet
In Sections 3.13.3, we explained the role of PointNet [34] in each of the three proposed models, namely Flow Matching PointNet, Diffusion PointNet, and the baseline PointNet. In this part, we describe the architecture of PointNet itself. In general, the pair $\mathcal{X}$ and $\mathcal{Y}$ denote the input and output (or the quantity to be predicted) of PointNet [34], respectively. In the earlier sections, we clarified what $\mathcal{X}$ and $\mathcal{Y}$ represent for each of the three models. In summary, in Flow Matching PointNet, $\mathcal{X}\in\mathbb{R}^{N\times(d+d_{\tau}+n_{\text{CFD}})}$ and $\mathcal{Y}\in\mathbb{R}^{N\times n_{\text{CFD}}}$; in Diffusion PointNet, $\mathcal{X}\in\mathbb{R}^{N\times(d+d_{\tau}+n_{\text{CFD}})}$ and $\mathcal{Y}\in\mathbb{R}^{N\times n_{\text{CFD}}}$; and in the baseline PointNet, $\mathcal{X}\in\mathbb{R}^{N\times d}$ and $\mathcal{Y}\in\mathbb{R}^{N\times n_{\text{CFD}}}$.
@@
The original PointNet model developed for computer graphics applications consists of two branches: a classification branch and a segmentation branch [34]. In this work, we employ the segmentation branch of PointNet [34]. Figure 2 illustrates the overall architecture of the segmentation branch of PointNet [34, 29]. At a high level, the architecture can be viewed as an encoder–decoder framework. A key advantage of PointNet [34] and its variants (see e.g., Ref. [35]) is their ability to process unordered sets of points that constitute a point cloud. In particular, PointNet is invariant under the $N!$ permutations of an input vector representing a point cloud with $N$ points. This permutation invariance is achieved through two main mechanisms. First, in both the encoder and decoder, each neural network layer uses a shared weight matrix and a shared bias vector across all neurons, forming what is known as a shared layer. Second, a symmetric function, one that is itself invariant to input ordering, is applied to encode global features from the input. As shown in Fig. 2, PointNet uses the max function as its symmetric function, as it has been shown to outperform alternatives such as the average function [34].
@@
In Fig. 2, the notation $(\mathcal{B}_{1},\mathcal{B}_{2})$ refers to two consecutive hidden layers with $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ neurons, whereas the triplet $(\mathcal{B}_{1},\mathcal{B}_{2},\mathcal{B}_{3})$ follows the same convention. As shown in Fig. 2, the initial feature extraction stage consists of two shared MLP blocks with layer sizes $(128,128)$ and $(128,256,2048)$. A symmetric max pooling operation is then applied to obtain a global feature vector of size $2048$. This global feature is concatenated with an intermediate tensor of size $N\times 128$, producing a combined representation of dimension $N\times 2176$. The subsequent processing stage employs two additional shared MLP blocks with sizes $(1024,512,256)$ and $(256,n_{\mathrm{CFD}})$. Consequently, the final output of PointNet is a tensor of size $N\times n_{\mathrm{CFD}}$, as depicted in Fig. 2. After each shared MLP layer, a rectified linear unit activation function,
@@
Further details about the architecture of PointNet can be found in Refs. [29, 30].
Table 2: Comparison between the performance of the Flow Matching PointNet, Diffusion PointNet, and PointNet models for prediction of the velocity and pressure fields for the test set containing 222 unseen geometries. Number of sampling for the Flow Matching PointNet and Diffusion PointNet is 100. In numbers reported in the form $\alpha_{1}\pm\alpha_{2}$, $\alpha_{1}$ denotes the mean error and $\alpha_{2}$ represents the corresponding standard deviation. Note that $\cdots||$ indicates the $L^{2}$ norm.
Flow Matching PointNet Diffusion PointNet PointNet
@@
(g) PointNet
(h) PointNet
(i) PointNet
Fig. 4: Comparison of histograms of the relative error in $L^{2}$ norm for the velocity and pressure fields predicted by the Flow Matching PointNet, Diffusion PointNet, and PointNet models.

### v2

@@ Abstract
We present two novel generative geometric deep learning frameworks, termed Flow Matching PointNet and Diffusion PointNet, for predicting fluid flow variables on irregular geometries by incorporating PointNet into flow matching and diffusion models, respectively. In these frameworks, a reverse generative process reconstructs physical fields from standard Gaussian noise conditioned on unseen geometries. The proposed approaches operate directly on point-cloud representations of computational domains (e.g., grid vertices of finite-volume meshes) and therefore avoid the limitations of pixelation used to project geometries onto uniform lattices, as is common in U-Net-based flow matching and diffusion models. In contrast to graph neural network–based diffusion models, Flow Matching PointNet and Diffusion PointNet do not exhibit high-frequency noise artifacts in the predicted fields. Moreover, unlike such approaches, which require auxiliary intermediate networks to condition geometry, the proposed frameworks rely solely on PointNet, resulting in a simple and unified architecture. The performance of the proposed frameworks is evaluated on steady incompressible flow past a cylinder, using a geometric dataset constructed by varying the cylinder’s cross-sectional shape and orientation across samples. The results demonstrate that Flow Matching PointNet and Diffusion PointNet achieve more accurate predictions of velocity and pressure fields, as well as lift and drag forces, and exhibit greater robustness to incomplete geometries compared to a vanilla PointNet with the same number of trainable parameters.
@@ 1 Introduction and motivation
Flow matching [1], diffusion models [2], and their advanced variants [3, 4, 5] are generative deep learning frameworks that have demonstrated success in computer graphics and computer vision, particularly for image and video generation [6], outperforming earlier generative approaches such as generative adversarial networks (GANs) [7, 8] and variational autoencoders (VAEs) [9]. At a high level, in supervised flow matching [1] and diffusion [2] frameworks, a forward process adds noise to the training data, while an inverse process predicts the clean fields. Several formulations of flow matching and diffusion models operate in a conditional setting, where generation is guided by auxiliary inputs such as textual prompts [4, 10] or low-resolution images [11, 12]. Motivated by their success in computer graphics, these models have been extended to computational physics, enabling a range of applications including predictions conditioned on material properties [13, 14, 15], high-fidelity field reconstruction from low-resolution data [16, 17], and prediction of physical fields conditioned on the geometry of the computational domain [13, 18, 19]. In this article, we focus on the latter case.
@@
Flow matching and diffusion models conditioned on the geometry of the computational domain fall within the category of geometric deep learning. These approaches are critical in computational physics, as they accelerate shape-optimization studies that require exploration of high-dimensional geometric parameter spaces [20]. Generative flow matching and diffusion models conditioned on geometry have been shown to outperform regression-based surrogate models for predicting physical fields (see e.g., Refs. [21, 22, 23]), such as convolutional neural networks (CNNs) [24], Fourier neural operators (FNOs) [25], and deep operator neural networks (DeepONets) [26], which directly map geometry to physical fields. In particular, traditional CNNs [24], FNOs [25], and DeepONets [26] produce a single output for a given geometry, whereas flow matching and diffusion models can generate multiple realizations for the same geometry, thereby quantifying model uncertainty. Moreover, diffusion models have demonstrated superior performance compared to GANs and VAEs in predicting physical field variables (see, e.g., Refs. [21, 27]). In geometric deep learning using flow matching and diffusion models, two steps are particularly important. First, how to represent irregular geometries; and second, how to condition flow matching and diffusion models on geometry from a computer science perspective.
@@
For the representation of irregular geometries, two common approaches are used. The first is a pixelation-based strategy, in which the geometry is projected onto uniform Cartesian grids. In this setting, Q. Liu and N. Thuerey [13] proposed a diffusion model conditioned on the geometry of airfoils to predict velocity and pressure fields. In their work [13], the airfoil geometry was represented using the pixelation approach, and a U-Net architecture [28, 2] was employed as the backbone neural network. The authors [13] performed geometric conditioning by concatenating the binary image of the airfoil with the velocity and pressure fields as the input to the U-Net, while noise was added only to the velocity and pressure fields during training (see Fig. 3 of Ref. [13]). A similar approach was taken by Zheng et al. [18], where the authors employed an encoder that maps a two-dimensional binary image of the geometry to a latent feature representation, which was then used in a diffusion transformer to condition the diffusion model on the geometry (see Fig. 1 of Ref. [18]). Similarly, J. Hu et al. [19] proposed a diffusion model conditioned on the geometry of obstacles to predict incompressible flow fields. The authors [19] conditioned the diffusion model on geometry by encoding the obstacle as a binary mask through a U-Net–style encoder and injecting this geometry representation into the diffusion process via cross-attention blocks (see Fig. 1 of Ref. [19]). Their results demonstrated that the proposed diffusion model outperformed CNNs and VAEs [19].
@@
The Flow Matching PointNet and Diffusion PointNet frameworks consist of a forward process and an inverse denoising process. In Diffusion PointNet, the forward process incrementally corrupts ground-truth velocity and pressure fields in fluid flows by adding standard Gaussian noise. PointNet is then trained to predict the noise added at each diffusion step [2]. This prediction is conditioned on the spatial coordinates of the point cloud, which represent the domain geometry, as well as on a sinusoidal time embedding [48]. The conditioning is implemented by concatenating the spatial-coordinate vectors and the sinusoidal-embedding vector with the corrupted velocity and pressure fields. In the inverse process, generation begins from samples drawn from a standard Gaussian noise distribution and is conditioned on the spatial coordinates of the point cloud and the sinusoidal embedding. During the inverse procedure, the trained PointNet predicts the noise at each stage and progressively reconstructs the velocity and pressure fields in a manner consistent with the imposed conditioning. In the Flow Matching PointNet model, PointNet is trained to predict a target vector field that corresponds to the temporal derivative of a continuously perturbed field with respect to an auxiliary time variable [1]. In the inverse process, this learned vector field is used to numerically integrate an ordinary differential equation, progressively reconstructing the velocity and pressure fields from an initial standard Gaussian noise distribution [1], using the same conditioning mechanism as in Diffusion PointNet. To investigate the performance of Flow Matching PointNet and Diffusion PointNet, the classical fluid dynamics problem of incompressible flow past a cylinder is considered, in which the cross-sectional shape of the cylinder varies across the dataset.
Note that combinations of diffusion models and point cloud neural networks in the computer graphics literature have been used in which, during the forward process, noise is added to the spatial coordinates of shapes and, during the inverse process, new three-dimensional (3D) geometries are generated [41, 49]. In contrast, in the present study, noise is added to and removed from the physical fields, while the spatial coordinates remain clean and unchanged. Additionally, Y. He et al. [50] proposed a framework for semantic segmentation of 3D point clouds using a diffusion model (see Fig. 1 of Ref. [50]). In their approach, the diffusion process is conditioned on two components: first, the geometry of the input represented by the point cloud, and second, the semantic context, which provides a prior estimate of the possible label types for the points in the point cloud. Although this framework is successful in computer graphics, it is not applicable to predicting physical quantities such as velocity and pressure fields, since these quantities are continuous rather than discrete integer labels, as in semantic segmentation.
@@ 2 Data preparation
Fig. 2: Architecture of Flow Matching PointNet, Diffusion PointNet, and Baseline PointNet. Labels in the form $(\mathcal{B}_{1},\mathcal{B}_{2})$ and $(\mathcal{B}_{1},\mathcal{B}_{2},\mathcal{B}_{3})$ are used to denote shared MLPs, as illustrated in Sect. 3.4. $n_{\mathrm{CFD}}$ denotes the number of flow variables and $N$ represents the number of points in the point clouds. In the Flow Matching PointNet architecture, the subscript $\tau$ is used to indicate $\boldsymbol{s}_{\tau}$ and $\boldsymbol{y}_{\tau}$ (see Sect. 3.1). In the Diffusion PointNet architecture, the subscript $t$ is used to show $\boldsymbol{s}_{t}$ and $\boldsymbol{y}_{t}$ (see Sect. 3.2). In Flow Matching PointNet and Diffusion PointNet, the outputs are, respectively, the prediction of the target field ($\boldsymbol{f}_{\text{target}}$) and the noisy field ($\boldsymbol{y}_{\text{noisy}}$), given the polluted velocity and pressure fields at steps $\tau$ and $t$, and conditioned on the geometry of the physical domain represented as a point cloud.
@@ 3.1 Generative Flow Matching PointNet
After training, Flow Matching PointNet generates the velocity and pressure fields conditioned on a geometry from the test set. Given $\boldsymbol{x}$ as the point cloud representing the target geometry from the test set, we generate three independent standard Gaussian fields representing $\boldsymbol{y}_{\text{noisy}}$ and integrate Eq. 12 from $\tau=1$ to $\tau=0$ using the explicit Euler method with $N_{\tau}$ steps, such that $\Delta\tau=1/N_{\tau}$. At each step,
@@ 3.2 Generative Diffusion PointNet
Similar to Sect. 3.1, let $\boldsymbol{y}_{\text{clean}}\in\mathbb{R}^{N\times n_{\text{CFD}}}$ represent the clean field, where $N$ and $n_{\text{CFD}}$ are defined as before, since our objective is to predict the two-dimensional velocity and pressure variables $u^{\prime}$, $v^{\prime}$, and $p^{\prime}$. Figure 3 illustrates the schematic of the forward and inverse processes in Diffusion PointNet. In the forward diffusion process, standard Gaussian noise is incrementally added to $\boldsymbol{y}_{\text{clean}}$ over $T$ discrete steps using a noise scheduler, producing a sequence of noisy fields $\{\boldsymbol{y}_{t}\}_{t=1}^{T}$, where $\boldsymbol{y}_{t}\in\mathbb{R}^{N\times n_{\text{CFD}}}$. In this study, we use a cosine noise scheduler defined as
@@
where $\epsilon_{u_{i}}$, $\epsilon_{v_{i}}$, and $\epsilon_{p_{i}}$ denote the ground-truth standard Gaussian noise, as similarly defined in Sect. 3.1, and $\tilde{\epsilon}_{u_{i}}$, $\tilde{\epsilon}_{v_{i}}$, and $\tilde{\epsilon}_{p_{i}}$ denote the corresponding noise components predicted by the Diffusion PointNet. Diffusion PointNet is optimized with the Adam algorithm [62] using a constant learning rate of 0.001 and a batch size of 256. For each training iteration, a random step $t$ is sampled independently for every point cloud in the training batch, and three independent standard Gaussian noises are generated, corresponding to the $u^{\prime}$, $v^{\prime}$, and $p^{\prime}$ variables, to construct the tensor $\boldsymbol{y}_{\text{noisy}}$ presented in Eq. 19. The corresponding noisy tensor $\boldsymbol{y}_{t}$ is then computed using Eq. 19.
Once training is complete, the Diffusion PointNet generates the corresponding velocity and pressure variables for a geometry selected from the test set by simulating the reverse diffusion process, as shown in Fig. 3. For a point cloud $\boldsymbol{x}$ representing a geometry from the test set, the reconstruction begins by generating three independent standard Gaussian noise fields to construct the tensor $\boldsymbol{y}_{T}$. The fields are then iteratively denoised from $t=T$ to $t=0$ using
@@ 3.3 Baseline PointNet
Now that the Flow Matching PointNet and Diffusion PointNet architectures have been described, the baseline PointNet can be introduced in a straightforward manner. The baseline PointNet for predicting fluid flow variables was first proposed by Kashefi et al. [29] and has since been adopted in several subsequent frameworks (see e.g., Refs. [64, 30, 65]). The version considered in this work is identical to that originally presented by Kashefi et al [29]. In the baseline PointNet, the input consists solely of the spatial coordinates of the point cloud representing the geometry of the domain, and the output corresponds to the physical fields of interest, as shown in the bottom panel of Fig. 2. Following the notation introduced in Sect. 3.1, the input is given by $\boldsymbol{x}\in\mathbb{R}^{N\times d}$, and consequently, the input tensor is $\mathcal{X}\in\mathbb{R}^{N\times d}$. The output consists of the velocity components and pressure field, forming an output tensor $\mathcal{Y}\in\mathbb{R}^{N\times n_{\mathrm{CFD}}}$. Hence, PointNet learns a direct mapping from $\mathcal{X}$ to $\mathcal{Y}$. The loss function is the mean squared error and can be written as
@@ 3.4 PointNet
In Sections 3.13.3, we explained the role of PointNet [34] in each of the three proposed models, namely Flow Matching PointNet, Diffusion PointNet, and the baseline PointNet. In this part, we describe the architecture of PointNet itself. In general, the pair $\mathcal{X}$ and $\mathcal{Y}$ denote the input and output (or the quantity to be predicted) of PointNet [34], respectively. In the earlier sections, we clarified what $\mathcal{X}$ and $\mathcal{Y}$ represent for each of the three models. In summary, in Flow Matching PointNet, $\mathcal{X}\in\mathbb{R}^{N\times(d+d_{\tau}+n_{\text{CFD}})}$ and $\mathcal{Y}\in\mathbb{R}^{N\times n_{\text{CFD}}}$; in Diffusion PointNet, $\mathcal{X}\in\mathbb{R}^{N\times(d+d_{t}+n_{\text{CFD}})}$ and $\mathcal{Y}\in\mathbb{R}^{N\times n_{\text{CFD}}}$; and in the baseline PointNet, $\mathcal{X}\in\mathbb{R}^{N\times d}$ and $\mathcal{Y}\in\mathbb{R}^{N\times n_{\text{CFD}}}$.
@@
The original PointNet model developed for computer graphics applications consists of two branches: a classification branch and a segmentation branch [34]. In this work, we employ the segmentation branch of PointNet [34]. Figure 2 illustrates the overall architecture of the segmentation branch of PointNet [34, 29]. At a high level, the architecture can be viewed as an encoder–decoder framework. A key advantage of PointNet [34] and its variants (see e.g., Ref. [35]) is their ability to process unordered sets of points that constitute a point cloud. In particular, PointNet is invariant under the $N!$ permutations of an input vector representing a point cloud with $N$ points [34]. This permutation invariance is achieved through two main mechanisms. First, in both the encoder and decoder, each neural network layer uses a shared weight matrix and a shared bias vector across all neurons, forming what is known as a shared layer. Second, a symmetric function, one that is itself invariant to input ordering, is applied to encode global features from the input. As shown in Fig. 2, PointNet uses the max function as its symmetric function, as it has been shown to outperform alternatives such as the average function [34].
@@
In Fig. 2, the notation $(\mathcal{B}_{1},\mathcal{B}_{2})$ refers to two consecutive hidden shared layers with $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ neurons, whereas the triplet $(\mathcal{B}_{1},\mathcal{B}_{2},\mathcal{B}_{3})$ follows the same convention. As shown in Fig. 2, the initial feature extraction stage consists of two shared MLP blocks with layer sizes $(128,128)$ and $(128,256,2048)$. A symmetric max pooling operation is then applied to obtain a global feature vector of size $2048$. This global feature is concatenated with an intermediate tensor of size $N\times 128$, producing a combined representation of dimension $N\times 2176$. The subsequent processing stage employs two additional shared MLP blocks with sizes $(1024,512,256)$ and $(256,n_{\mathrm{CFD}})$. Consequently, the final output of PointNet is a tensor of size $N\times n_{\mathrm{CFD}}$, as depicted in Fig. 2. After each shared MLP layer, a rectified linear unit activation function,
@@
Further details about the architecture of PointNet can be found in Refs. [34, 29, 30].
Table 2: Comparison between the performance of the Flow Matching PointNet, Diffusion PointNet, and baseline PointNet models for prediction of the velocity and pressure fields for the test set containing 222 unseen geometries. Number of sampling for the Flow Matching PointNet and Diffusion PointNet is 100. In numbers reported in the form $\alpha_{1}\pm\alpha_{2}$, $\alpha_{1}$ denotes the mean error and $\alpha_{2}$ represents the corresponding standard deviation. Note that $\cdots||$ indicates the $L^{2}$ norm.
Flow Matching PointNet Diffusion PointNet Baseline PointNet
@@
(g) Baseline PointNet
(h) Baseline PointNet
(i) Baseline PointNet
Fig. 4: Comparison of histograms of the relative pointwise error in $L^{2}$ norm for the velocity and pressure fields predicted by the Flow Matching PointNet, Diffusion PointNet, and baseline PointNet models.